a rope of length 44cm is used to form a rectangle.if the length of rectangle is six more than the width,find the length of the rectangle
Answer by Cromlix(4381)

Hi there,
Using a rope to make a rectangle,
the rope is acting as a perimeter.
Make width = x
Length = 6 + x
Perimeter = 2(Width) + 2(Length)
44 = 2(x) + 2(6 + x)
44 = 2x + 12 + 2x
Collect like terms:
2x + 2x = 44 - 12
4x = 32
x = 8
Width 8 cm
Length = 14 cm
Hope this helps :-)